OpenStudy (kpop4life123):

What is the volume of a cone with a diameter of 18 cm and a height of 10 cm? Use 3.14 for pi. Round your answer to the nearest tenth

pooja195 (pooja195):

\[\huge~\rm~\bf~V=\pi~r^2\frac{ h }{ 3}\]

pooja195 (pooja195):

To find the radius divide the diameter by 2 18/2=?

OpenStudy (kpop4life123):

9

pooja195 (pooja195):

Good \[\huge~\rm~\bf~V=(3.14)(3^2)(\frac{ 10 }{ 3})\]

pooja195 (pooja195):

3^2=?

OpenStudy (kpop4life123):

9

pooja195 (pooja195):

Good Now just solve for V Good \[\huge~\rm~\bf~V=(3.14)(9)(\frac{ 10 }{ 3})\]

OpenStudy (kpop4life123):

dont we have to divide 10 by 3?

pooja195 (pooja195):

Yeah

pooja195 (pooja195):

I suggest using a calculator for this part

OpenStudy (kpop4life123):

wait let me see

OpenStudy (kpop4life123):

94.2? @pooja195

pooja195 (pooja195):

That's correct

OpenStudy (kpop4life123):

thank you :)

pooja195 (pooja195):

yw ^_^

pooja195 (pooja195):

Wait hold up

pooja195 (pooja195):

I messed up the radius is 9 not 3 >.< \[\huge~\rm~\bf~V=(3.14)(9^2)(\frac{ 10 }{ 3})\] \[\huge~\rm~\bf~V=(3.14)(81)(\frac{ 10 }{ 3})\] Now solve for V

OpenStudy (kpop4life123):

847.8?

pooja195 (pooja195):

There we go ok yes thats correct

OpenStudy (kpop4life123):

can you help me on one more:)

pooja195 (pooja195):

ok :) post it ^_^

OpenStudy (kpop4life123):

⦁ A cone has a radius of 6 inches and a height of 11 inches. What is the volume of the cone?

pooja195 (pooja195):

\[\huge~\rm~\bf~V=\pi~r^2\frac{ h }{ 3}\] Same formula \[\huge~\rm~\bf~V=(3.14)~(6)^2(\frac{ 11 }{ 3})\] solve for v

OpenStudy (kpop4life123):

414.48?

pooja195 (pooja195):

i got 414.69 but it doesnt really matter :P

pooja195 (pooja195):

Well done =)

OpenStudy (kpop4life123):

:)